# Systematic derivation of perturbative unitarity bounds for any models

### v1

@@ I Introduction
An important illustration of these theoretical constraints is the requirement of the unitarity of the scattering matrix or S-matrix, which leads to the so-called perturbative unitarity bounds. The foundational work of Lee, Quigg, and Thacker [3] demonstrated its power by deriving an upper bound on the SM Higgs mass. Their analysis leveraged the Goldstone boson equivalence theorem, which states that at high energies, the scattering amplitudes of longitudinal gauge bosons are equivalent to their corresponding Goldstone bosons. This allows unitarity constraints to be efficiently applied to the scalar sector of the theory. The conventional approach, widely used in studies of BSM models, is to consider the high-energy limit where the center-of-mass energy far exceeds all particle masses. In this regime, the calculation simplifies as diagrams with propagators are suppressed, leaving only contact interactions.
In this letter we present a general method, which systematically derives the $2\rightarrow 2$ scattering matrix in the high-energy limit for models with arbitrary gauge groups and field representations. We can then compute the eigenvalues of the S-matrix and apply the perturbative unitarity bounds on them. This method is given in our Mathematica package anyPUB. This package, accompanied with an example file, can be found at https://github.com/nicobenincasa/anyPUB. We then apply our framework to the minimal left-right symmetric model (MLRSM), deriving explicit constraints and comparing our results with the existing literature. To further demonstrate the generality of our formalism, we employ it in the context of the Pati-Salam model, providing the first complete calculation of its perturbative unitarity constraints.
@@ IV Discussion
We claim that anyPUB works for any models. We tested it for a range of models such as SM + real/complex singlet [4, 5], $\mathbb{Z}_{2}$-symmetric 2HDM [6], 2HD+a model [7], IDM [8], dark $SU(2)$ model [9], $\mathbb{Z}_{3}$-2HDM + complex singlet [10], $U(1)\times U(1)$-symmetric 3HDM [11], Higgs triplet model [12], economical 331 model [13], Georgi-Machacek model [14], type-II seesaw model [15], minimal left-right symmetric model [16], where we could successfully obtain the analyical results found in the literature. However for models like general 2HDM [17], $\mathbb{Z}_{3}$-symmetric 3HDM [11] or Pati-Salam model [18], the size of some matrix blocks were too large to compute the corresponding eigenvalues analytically.
Actually for $SU(2)\otimes U(1)$ gauge theories with an arbitrary number of scalar fields, either singlets or doublets of $SU(2)$, a recent Mathematica package BounDS can systematically obtain a basis in terms of quantum numbers $|Q,Y,T\rangle$, thereby blocks of the scattering matrix are naturally smaller and it is then easier to extract the eigenvalues analytically [19]. While our method also works for such models, it becomes trickier to obtain all eigenvalues analytically for more complex versions of these models like the general NHDM. However, we can still use our method to obtain the eigenvalues numerically. Moreover, unlike BounDS, it also works for models involving any gauge groups and multiplet representations, e.g the minimal left-right symmetric model consisting in bi-doublets and triplets (see Section IV.1). Finally, the function MakeBlockDiagonal that block-diagonalises matrices can also be used for large scattering matrices obtained with the SARAH Mathematica package [20], in order to compute the eigenvalues more easily. Let us now illustrate our method with two specific models: the MLRSM and the Pati-Salam model.
@@ IV.1 Minimal left-right symmetric model
$\Delta_{L,R}=\begin{pmatrix}\delta^{+}_{L,R}/\sqrt{2}&\delta^{++}_{L,R}\\ \delta^{0}_{L,R}&-\delta^{+}_{L,R}/\sqrt{2}\end{pmatrix}$. (6)
@@
$(20\lambda_{1}^{2}+240\lambda_{1}\lambda_{2}+88\lambda_{1}\lambda_{3}+96\lambda_{1}\lambda_{5}+48\lambda_{1}\lambda_{6}$
@@ IV.2 Pati-Salam model
$+\lambda_{R\Phi 3}{\rm{Tr}}\left[\phi_{R}\phi_{R}^{\dagger}\Phi^{\dagger}\Phi\right].$ (24)
@@ V Conclusions
We have derived the perturbative unitarity bounds in the case of the minimal left-right symmetric model and showed that our results are very different than those of [16], highlighting the sources of discrepancies. We obtained only 21 independent eigenvalues compared to 64 in their paper. Moreover, the expressions of our eigenvalues are much simpler than theirs. Next, we derived for the first time the perturbative unitarity bounds in the Pati-Salam model. We could obtain 10 eigenvalues analytically, while the 13 remaining ones should be computed numerically.
@@ Appendix A How to use anyPUB?
A.1 Example 1: Standard Model + real singlet with $\mathbb{Z}_{2}$ symmetry
@@
In order to block diagonalise the scattering matrix, convert it into an adjacency matrix by putting all non-zero off-diagonal elements of the S-matrix to 1, and the remaining elements to zero [23]:
@@
[…]
The adjacency matrix $A$ of a graph should be interpreted as the following: an non-zero element $A_{ij}$ corresponds to an oriented edge $i\rightarrow j$ in the graph, from the vertex labeled by $i$ to the vertex labeled by $j$. The scattering matrix is hermitian at tree-level, thus the resulting adjacency matrix is symmetric, leading to an undirected graph (for the vertex $i$ and $j$, one both has the edge $i\rightarrow j$ and the edge $j\rightarrow i$). Therefore we only focus on the upper triangle. The non zero-elements are at the following positions:
@@ 3 VSinglet, assumption]["AllDistinctEigenvalues"]]
A.2 Example 2: Two-Higgs-doublet model
@@
A.2.1 $\lambda_{6}\neq 0,\lambda_{7}=0$ and $\lambda_{6}\neq 0,\lambda_{7}\neq 0$

### v2

@@ I Introduction
An important illustration of these theoretical constraints is the requirement of the unitarity of the scattering matrix or S-matrix, which leads to the so-called perturbative unitarity bounds. The foundational work of Lee, Quigg, and Thacker [22] demonstrated its power by deriving an upper bound on the SM Higgs mass. Their analysis leveraged the Goldstone boson equivalence theorem, which states that at high energies, the scattering amplitudes of longitudinal gauge bosons are equivalent to their corresponding Goldstone bosons. This allows unitarity constraints to be efficiently applied to the scalar sector of the theory. Note that perturbative unitarity constraints on generic Yukawa and vector interactions have been analysed in [3, 7]. Regarding the scalar sector, the conventional approach, widely used in studies of BSM models, is to consider the high-energy limit where the center-of-mass energy far exceeds all particle masses. In this regime, the calculation simplifies as diagrams with propagators are suppressed, leaving only contact interactions.
In this letter we present a general method, which systematically derives the $2\rightarrow 2$ scattering matrix in the high-energy limit for models with arbitrary gauge groups and field representations. We can then compute the eigenvalues of the S-matrix and apply the perturbative unitarity bounds on them. This method is given in our Mathematica package anyPUB. This package, accompanied with an example file, can be found at https://github.com/nicobenincasa/anyPUB. We then apply our framework to the minimal left-right symmetric model (MLRSM), deriving explicit constraints and comparing our results with the existing literature, showing that the previously obtained results are wrong. To further demonstrate the generality of our formalism, we employ it in the context of the Pati-Salam model, providing the first complete calculation of its perturbative unitarity constraints.
@@ IV Discussion
We claim that anyPUB works for any models. We tested it for a range of models such as SM + real/complex singlet [19, 2], $\mathbb{Z}_{2}$-symmetric 2HDM [14], 2HD+a model [5], IDM [8], dark $SU(2)$ model [9], $\mathbb{Z}_{3}$-2HDM + complex singlet [10], $U(1)\times U(1)$-symmetric 3HDM [11], Higgs triplet model [21], economical 331 model [20], Georgi-Machacek model [4] and type-II seesaw model [6], where we could successfully obtain the analyical results found in the literature. However for models like general 2HDM [15] or $\mathbb{Z}_{3}$-symmetric 3HDM [11], the size of some matrix blocks were too large to compute the corresponding eigenvalues analytically.
Actually for $SU(2)\otimes U(1)$ gauge theories with an arbitrary number of scalar fields, either singlets or doublets of $SU(2)$, a recent Mathematica package BounDS can systematically obtain a basis in terms of quantum numbers $|Q,Y,T\rangle$, thereby blocks of the scattering matrix are naturally smaller and it is then easier to extract the eigenvalues analytically [23]. While our method also works for such models, it becomes trickier to obtain all eigenvalues analytically for more complex versions of these models like the general NHDM. However, we can still use our method to obtain the eigenvalues numerically. Moreover, unlike BounDS, it also works for models involving any gauge groups and multiplet representations, e.g the minimal left-right symmetric model consisting in bi-doublets and triplets (see Section IV.1).
Alternatively, one can use the SARAH Mathematica package [17], which is designed to compute a variety of theoretical quantities. In addition to perturbative unitarity analysis, SARAH can produce outputs such as mass matrices, renormalization group equations, and model files for external tools, enabling further phenomenological studies in combination with dedicated software. While SARAH is quite efficient when applied to pre-existing models, the construction and implementation of a new model demand substantial effort. Next, regarding perturbative unitarity, SARAH considers quantum numbers and conservation laws such that only states that are allowed to mix are considered when deriving the $2\rightarrow 2$ S-matrix. However, for models with large representations and many degrees of freedom, the process is computationally intensive and, in practice, unfeasible. Finally, note that the anyPUB function MakeBlockDiagonal that block-diagonalises matrices can also be used for large scattering matrices obtained with SARAH, in order to compute the eigenvalues more easily. Let us now illustrate our method with two specific models: the MLRSM and the Pati-Salam model.
@@ IV.1 Minimal left-right symmetric model
$\Delta_{L,R}=\begin{pmatrix}\delta^{+}_{L,R}/\sqrt{2}&\delta^{++}_{L,R}\\ \delta^{0}_{L,R}&-\delta^{+}_{L,R}/\sqrt{2}\end{pmatrix}$, (6)
with quantum numbers given in Table 1.
Fields SU(3)C SU(2)L SU(2)R U(1)B-L
$\Phi$ 1 2 2 0
$\Delta_{L}$ 1 3 1 2
$\Delta_{R}$ 1 1 3 2
Table 1: Scalar fields and their quantum numbers in the minimal left-right symmetric model.
@@
$+(20\lambda_{1}^{2}+240\lambda_{1}\lambda_{2}+88\lambda_{1}\lambda_{3}+96\lambda_{1}\lambda_{5}+48\lambda_{1}\lambda_{6}$
@@
Finally, let us show how these perturbative unitarity bounds constrain the parameter space in the MLRSM. In this model, the mass of the heavy scalar states is made of one leading-order contribution proportional to the vacuum expectation value (VEV) $v_{R}$ from $\langle\Delta_{R}\rangle$, while the sub-leading contribution is proportional to the electroweak VEV $v_{\text{EW}}$, with $v_{\text{EW}}\ll v_{R}$. The $v_{R}$ part of these masses is proportional to a linear combination of $\lambda_{5},\lambda_{6},\lambda_{7},\lambda_{12}$, while $\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{8},\lambda_{9},\lambda_{10},\lambda_{11}$ only contribute to the subdominant part proportional to $v_{\text{EW}}$ [24] (note that $\lambda_{12}$ also contributes to it [13]). Given that this model contains 12 quartic couplings, we assume a mass degeneracy at leading-order for the heavy scalars, thus implying $\lambda_{M}=\lambda_{5}=\lambda_{6}=(\lambda_{7}-2\lambda_{5})/4=\lambda_{12}/4$. Following [24], we also consider $\lambda_{u}=\lambda_{2}=\lambda_{3}=\lambda_{4}=\lambda_{8}=\lambda_{9}=\lambda_{10}=\lambda_{11}$. In Figure 1, we show the region of the parameter space, in the plane $\lambda_{u}-\lambda_{M}$, which satisfies the perturbative unitarity bounds in the MLRSM. We can see that these bounds constrain the quartic couplings to be small: $|\lambda_{u}|<0.7$ and $|\lambda_{M}|<0.8$, thus implying $M\lesssim 1.25~v_{R}$, with $M$ the mass of the heavy scalar states.
Figure 1: Region of the parameter space allowed by perturbative unitarity bounds in the MLRSM.
@@ IV.2 Pati-Salam model
$+\lambda_{R\Phi 3}{\rm{Tr}}\left[\phi_{R}\phi_{R}^{\dagger}\Phi^{\dagger}\Phi\right],$ (24)
where quantum numbers of the fields are given in Table 2.
Fields SU(4) SU(2)L SU(2)R
$\phi_{R}$ 4 1 2
$\Phi$ 1 2 2
Table 2: Scalar fields and their quantum numbers in the Pati-Salam model.
@@
Finally, let us show how these perturbative unitarity bounds constrain the parameter space in the Pati-Salam model. In this model, the mass of the heavy scalar states is made of one leading-order contribution proportional to the vacuum expectation value $v_{R}$ from $\langle\phi_{R}\rangle$, while the sub-leading contribution is proportional to the electroweak VEV $v_{\text{EW}}$, with $v_{\text{EW}}\ll v_{R}$. The $v_{R}$ part of these masses is proportional to a linear combination of $\lambda_{R1},\lambda_{R2},\lambda_{R\phi 3}$, while $\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{R\phi 1},\lambda_{R\phi 2}$ only contribute to the subdominant part proportional to $v_{\text{EW}}$. As in the previous section we then assume a mass degeneracy at leading-order for the heavy scalars, thus implying $\lambda_{M}=\lambda_{R1}=-3\lambda_{R2}/2=-(\lambda_{R\phi 3}+4\lambda_{R2})/4$. Next we also consider $\lambda_{u}=\lambda_{2}=\lambda_{3}=\lambda_{4}=\lambda_{R\phi 1}=\lambda_{R\phi 2}$. In Figure 2, we show the region of the parameter space, in the plane $\lambda_{u}-\lambda_{M}$, which satisfies the perturbative unitarity bounds in the Pati-Salam model. We can see that these bounds constrain the quartic couplings to be small, although the bounds are looser than in the MLRSM: $|\lambda_{u}|<1$ and $|\lambda_{M}|<2.4$, thus implying $M\lesssim 2.2~v_{R}$, with $M$ the mass of the heavy scalar states.
Figure 2: Region of the parameter space allowed by perturbative unitarity bounds in the Pati-Salam model.
@@ V Conclusions
We have derived the perturbative unitarity bounds in the case of the minimal left-right symmetric model and shown that our results are very different than those of [24], highlighting the sources of discrepancies. We obtained only 21 independent eigenvalues compared to 64 in their paper. Moreover, the expressions of our eigenvalues are much simpler than theirs. Next, we derived for the first time the perturbative unitarity bounds in the Pati-Salam model. We could obtain 10 eigenvalues analytically, while the 13 remaining ones should be computed numerically. Finally, for both models, we have graphically shown how these perturbative unitarity bounds constrain the parameter space.
@@ Appendix A How to use anyPUB?
A.1 Definition of functions
In this section, we provide a description for the functions used in anyPUB:
  • ScatteringMatrix[quarticPotential, assump- tion], with quarticPotential the quartic part of the scalar potential and assumption specifying which quartic couplings are real, returns an association between keys and values: FieldList is the list of independent real scalar fields in quarticPotential, PairsList is the list of all unordered pairs of these fields, ScatteringPairs is a nested list of all possible unordered combinations of these pairs, Vertices is a nested list computed four-point vertices for each scattering pair and ScatteringMatrix is the S-matrix.
  • MakeBlockDiagonal[matrix], with matrix the matrix to diagonalise, returns an association between keys and values: AdjacencyMatrix is the adjacency matrix obtained from matrix, Graph is the graph associated to this adjacency matrix, Components is the list of rows/columns index used for each connected subgraphs of the graph and BlockMatrix is the block-diagonalised version of matrix.
  • EigenvaluesScatteringMatrix[quarticPotent- ial, assumption], with quarticPotential the quartic part of the scalar potential and assumption specifying which quartic couplings are real, returns an association between keys and values: EigenvaluesPerBlock is the list of eigenvalues per irreducible block of the block-diagonalised matrix, while AllDistinctEigenvalues provide the list of distinct eigenvalues of the S-matrix.
  • MatrixMinimalPolynomial[matrix, x, number- DistinctEigenvalues] returns the minimal polynomial of matrix using x as a variable. The number of distinct eigenvalues, number- DistinctEigenvalues is optional, it is set to 0 by default. If it is larger than 0, then this function avoids computing the null space of matrix as long as the number of iterations is smaller than or equal to numberDistinctEigenvalues.
  • ReducedCharacteristicPolynomial[character- isticPolynomial, eigenvaluesList, x] returns the polynomial quotient of character- isticPolynomial in terms of x, by using a set of known eigenvalues, eigenvaluesList.
  • ReducedPolynomialNewton[matrix, eigenvalu- esList, multiplicityList, x] returns a polynomial in terms of x, via the Newton’s identities, by using a set a known eigenvalues, eigenvaluesList, and their corresponding multiplicity, multiplicityList.
A.2 Example 1: Standard Model + real singlet with $\mathbb{Z}_{2}$ symmetry
@@
In order to block diagonalise the scattering matrix, convert it into an adjacency matrix by putting all non-zero off-diagonal elements of the S-matrix to 1, and the remaining elements to zero [18]. This adjacency matrix can be obtained via
@@
The adjacency matrix $A$ of a graph should be interpreted as the following: an non-zero element $A_{ij}$ corresponds to an oriented edge $i\rightarrow j$ in the graph, from the vertex labeled by $i$ to the vertex labeled by $j$. The scattering matrix is hermitian at tree-level, thus the resulting adjacency matrix is symmetric, leading to an undirected graph (for the vertex $i$ and $j$, one both has the edge $i\rightarrow j$ and the edge $j\rightarrow i$). Therefore we only focus on the upper triangle. The non zero-elements of the adjacency matrix are at the following positions:
@@ 3 VSinglet, assumption]["AllDistinctEigenvalues"]]
A.3 Example 2: Two-Higgs-doublet model
@@
A.3.1 $\lambda_{6}\neq 0,\lambda_{7}=0$ and $\lambda_{6}\neq 0,\lambda_{7}\neq 0$